# Variational Shape Inference for Grasp Diffusion on

### v1

@@ I Introduction
Given an object’s geometry, there are typically multiple feasible grasp solutions; hence, it is of interest to encode the multimodality of the task and capture the entire solution set, which can be fine-tuned or sub-sampled based on task-specific criteria [7, 8]. Denoising diffusion models [9, 10] have gained significant popularity across computer vision [11, 12, 13], language modeling [14, 15, 16], robotics [17, 18, 19], and beyond, due to their ability to capture highly complex multimodal distributions without mode collapse. In this work, we leverage diffusion models to learn the distribution of stable grasps conditioned on object geometry.
@@
To enable a robust, generalizable, and multimodal grasp synthesis framework, we propose an approach centered around robust geometry inference. Our pipeline uses learned geometric features for stable, multimodal grasp synthesis through a two-stage process. We first learn a variational shape encoding guided by INRs, which maps object point clouds to a shape-latent space that enables robust shape inference using imperfect observations. We then use the learned shape features for generating grasp poses via diffusion on the manifold. Similar to Urain et al. [8], our framework can be directly incorporated into multi-objective optimization, allowing us to tune the generated grasp poses to include feasibility constraints. We demonstrate this capability by introducing a test-time optimization routine that uses a smooth, differentiable representation of the robot gripper to fine-tune the generated grasp poses for improved stability and success rates. We present results on noisy and sparse point clouds, and demonstrate the effectiveness of our method in real-world settings.
@@ II-A Multimodal Grasp Synthesis
In recent years, several methods have been proposed for generating a diverse set of grasps conditioned on object geometry. Mousavian et al. [7] propose a conditional variational autoencoder (cVAE) that learns to compress and reconstruct grasps conditioned on object point clouds. However, VAEs are known to exhibit mode collapse and failure in fully capturing the data distribution. To improve performance, this work couples the grasp generator with a grasp classifier, which is expensive to train.
Diffusion models have demonstrated improved multimodal distribution modeling and stable learning dynamics while avoiding mode collapse [27], leading to their emergence in robotics applications [17, 18]. Urain et al. [8]’s formulation is closest to ours, where a diffusion model is learned via denoising score matching guided by an object reconstruction cost, which represents the current state-of-the-art in multimodal grasp synthesis. Our framework better models geometric structure in the presence of noise and calibration errors by learning to project geometric observations to a latent space of shapes and using it to infer shape features to guide grasp synthesis.
Several recent works have extended diffusion-based grasp synthesis in various directions. Chen et al. [28] propose using stochastic interpolants to sample from an informed prior distribution instead of a Gaussian and bridge to the target distribution. Singh et al. [29] use a part-guided diffusion approach to generate grasps tailored to areas of interest instead of the whole object. Barad et al. [30] propose learning a latent space where diffusion on grasps is performed, and use language to condition on grasp regions. Carvalho et al. [31] propose a denoising diffusion probabilistic model for grasp synthesis, and optimize the grasps at test-time using collision spheres. Similarly, we propose a test-time grasp pose optimization technique that boosts performance without additional training on expert grasps. Song et al. [32] present an approach combining dense prediction with multimodal grasp generation via diffusion models. Lim et al. [33] proposes a novel equivariant lifting layer for constructing -equivariant time-dependent velocity fields in flow matching. However, it performs worse in our experience, possibly requiring a significantly longer training time to handle the set of challenging objects used in our setup.
@@ III-B Signed Distance Fields
Signed Distance Fields (SDFs) implicitly represent geometry by providing, at every point in space, the distance to the nearest surface, with a sign denoting the interior (negative) or exterior (positive). Formally, we let denote the signed distance field. The surface is then represented as the zero level-set of its SDF: . SDFs are differentiable almost everywhere, and where the gradient exists, its negative points toward the nearest surface. Additionally, the gradient satisfies the Eikonal equation: . In recent years, much work has focused on implicitly representing 3D geometries via neural SDFs [25, 21], where the SDF is parameterized by a neural network. This allows querying at arbitrary resolutions with a small memory footprint and readily yields gradients via automatic differentiation. In this work, we use signed distance fields to learn shape features, which we transfer to stable grasp synthesis. We also use a learned SDF of the robot gripper for test-time pose optimization of the gripper given object point cloud.
@@ IV-A Shape Inference via Variational Autoencoding
High-quality grasp synthesis requires understanding geometric features that can transfer across object shapes and be robust to imperfect observations. While inferring shape features from point clouds is an extensively researched task [46, 39, 42], multimodal frameworks that learn shape priors in a latent space offer superior robustness to observation noise and capture a broader space of plausible shape hypotheses [47, 48]. Following Chou et al. [42], we learn SDF representations from object point clouds using a PointNet-based variational autoencoder (PointVAE) architecture that encodes useful geometric priors.
@@ IV-C Test-time Pose Optimization
where balance the influence of the objective functions with the score function. In our experiments, we use . We employ the SIREN architecture [20] to model due to its infinite differentiability and proven effectiveness for smooth, differentiable objectives [26].
@@ V Experiments
We evaluate our approach against state-of-the-art multimodal grasp synthesis frameworks: cVAE [7], -Diff [8], Bridger [28], and GLDM [30]. All methods are trained on an NVIDIA A100 (40 GB) GPU for hours using ten object categories from the ACRONYM dataset [56]: Book, Bottle, Bowl, Cap, CellPhone, Cup, Hammer, Mug, Scissors, and Shampoo. We implement our method in PyTorch [57] and evaluate grasp success and stability in IsaacGym [58]. Additionally, we demonstrate zero-shot sim-to-real transfer on a pick-and-place task without additional training.
@@ V-B Grasp Synthesis
| Methods | Performance Metrics | ||
| S.R. | EMD | Time (s) | |
| cVAE [7] | |||
| SE(3)-Diff [8] | |||
| Bridger [28] | |||
| GLDM [30] | |||
| Ours | |||
| Ours + GrOpt | |||
| w/VNN | |||
| w/VNN+ | |||
@@
We evaluate grasp synthesis performance on 52 unseen object instances from the ACRONYM dataset, measuring average success rate (SR) and an empirical Earth Mover’s Distance (EMD) with respect to the ground truth distribution as indicators of grasp quality and diversity, respectively. Tbl.˜II presents our main results, with corresponding distributions shown in Fig.˜4.
@@
The cVAE baseline suffers due to training on a limited set of objects for grasp synthesis, while also demonstrating a lack of diversity due to mode collapse [8]. Furthermore, to boost its performance the cVAE depends on guidance from a grasp classifier [7] which is expensive to train. GLDM’s emphasis on learning a latent grasp space, which we hypothesize is not needed due to already working in a low dimension, proves less effective. Although the SMLD formulation of -Diff achieves reasonable performance, we believe it is limited by the VNN-based shape encoder. Bridger improves on -Diff by employing stochastic interpolants, achieving a higher SR with significantly lower inference time. Our method achieves the highest success rate of , outperforming the previous state-of-the-art by , while maintaining competitive diversity (EMD = ) and reasonable inference time ( seconds).
@@
Key to our performance is an improved grasp synthesis architecture, a robust latent space for shape feature learning, and pre-training the shape backbone for object reconstruction. To see that the improved grasp synthesis architecture alone does not contribute to this performance improvement, we train "w/VNN" which yields lower performance. We also show the usefulness of pre-training for learning shape features by training "w/VNN+", where the VNN is pretrained via a surface reconstruction objective (described in Sec.˜V-A). Doing so boosts grasp synthesis performance compared to "w/VNN" but underperforms compared to Ours. This observation is supported by the shape inference results in Sec.˜V-A. Finally, our proposed test-time pose optimization method (Ours + GrOpt) boosts success rate to by encoding collision avoidance and pinch stability objectives. While this optimization slightly reduces diversity due to local refinement, it improves grasp quality and reduces variance. Qualitative results in Fig.˜5 illustrate the higher grasp quality achieved by our method across a diverse set of objects.
@@
V-C Point Cloud Sparsity
Real-world perception systems often produce sparse or noisy point clouds due to sensor limitations and environmental factors. We evaluate robustness by testing on point clouds of varying density, from 128 to 2048 points, where the training default is 1024. We exclude GLDM in this analysis as it does not ingest point clouds of varying sizes. Fig.˜6 shows that methods using standard point cloud encoders experience significant performance degradation as point cloud sparsity increases. In contrast, our approach has consistent performance across all sparsity levels, with success rates remaining above even with highly sparse inputs. This robustness stems from our variational shape inference, which learns a structured latent space that captures geometric priors and compensates for sparsity. This emphasizes the case for using the PointVAE architecture as a backbone for shape feature learning.
V-D Real-world Deployment
To validate the effectiveness of our variational shape inference approach with real sensor data, we deploy our method on a Franka Research 3 robot arm equipped with a Panda Hand gripper in a pick-and-place task. The experimental setup uses a Realsense D435i camera to capture the robot’s workspace, with segmented object point clouds extracted using SAM [61] as shown in Fig.˜7. Collision-free trajectories to and from the generated grasp poses are computed using the RRT-Connect implementation from VAMP [62]. Video demonstrations are provided in the supplementary material.
All methods are deployed without additional training or fine-tuning, evaluating zero-shot sim-to-real transfer capabilities. As presented in Tbl.˜III, our method achieves consistently high performance across all test objects, with particularly robust results for challenging geometries such as the Jug and Klein Bottle (K-Bottle). The Jug is significantly larger compared to other objects, and the Klein Bottle features smoothly varying curvature and complex geometry. These challenging objects are further complicated by real-world factors such as sensor noise and point cloud registration errors. Such conditions create substantial difficulties for baseline methods that rely on standard point cloud encoders, while highlighting the advantages of our variational shape inference approach (trained only on noise-free point clouds), which maintains robust performance despite measurement error. This robustness directly validates our core hypothesis that learning structured geometric priors through variational autoencoding improves grasp synthesis.
VI Conclusion
In this work, we bring together approaches from generative modeling and implicit neural representations to enable robust and generalizable grasp synthesis. We employ a shape inference formulation parameterizing SDFs conditioned on object point clouds, and use the learned model to formulate grasp synthesis via a denoising diffusion model. Our proposal achieves superior performance compared to the state of the art in multimodal grasp synthesis, and demonstrates robustness in the presence of observation noise. Additionally, our test-time optimization technique further improves grasp quality without requiring additional training from expert data.
This work furthers the case for the usefulness of transferring approaches for geometry understanding to grasp synthesis [6]. The demonstrated robustness to point cloud sparsity and the successful real-world experiments underscore the practical value of variational shape inference for robotic manipulation. Therefore, in the future, large pretraining for shape inference may immediately be made useful for grasp synthesis in a few-shot manner [63].

### v2

@@ I Introduction
Given an object’s geometry, there are typically multiple feasible grasp solutions; hence, it is of interest to encode the multimodality of the task and capture the entire solution set, which can then be sampled based on task-specific criteria [7, 8]. Denoising diffusion models [9, 10] have gained significant popularity across computer vision [11, 12, 13], language modeling [14, 15, 16], robotics [17, 18, 19], and beyond, due to their ability to capture highly complex multimodal distributions without mode collapse, stable training [20], and flexibility in guiding the sampling process at test-time [11]. In this work, we leverage diffusion models to learn the distribution of stable grasps conditioned on object geometry.
@@
To enable a robust, generalizable, and multimodal grasp synthesis framework, we propose an approach centered around robust geometry inference. Our pipeline uses learned geometric features for stable, multimodal grasp synthesis through a two-stage process. We first learn a variational shape encoding guided by INRs, which maps object point clouds to a shape-latent space that enables robust shape inference using imperfect observations. We then use the learned shape features for generating grasp poses via diffusion on the manifold. Similar to Urain et al. [8], our framework can be directly incorporated into multi-objective optimization, allowing us to tune the generated grasp poses to include feasibility constraints. We demonstrate this capability with a test-time optimization routine that uses a smooth, differentiable representation of the robot gripper to fine-tune the generated grasp poses for improved success rates. We present results on noisy, sparse, and partial point clouds, and demonstrate the effectiveness of our method in real-world settings. We release our code as open source111https://github.com/stalhabukhari/vsigd.
@@ II-A Multimodal Grasp Synthesis
Many methods have been proposed for generating a diverse set of grasps conditioned on object geometry. Mousavian et al. [7] propose a conditional variational autoencoder (cVAE) that learns to reconstruct grasps conditioned on object point clouds. However, VAEs are known to exhibit mode collapse and fail to fully capture the data distribution.
Diffusion models demonstrate improved multimodal distribution modeling and stable learning dynamics, avoiding mode collapse [20], and controllable sampling trade-offs and guidance mechanisms [11], leading to their use in robotics [17, 18]. Urain et al. [8]’s formulation is closest to ours, where a diffusion model is learned via denoising score matching guided by an object reconstruction cost, which represents the current state-of-the-art in multimodal grasp synthesis. Our framework better models geometric structure in the presence of noise and calibration errors by learning to project geometric observations to a latent space of shapes and using it to infer shape features to guide grasp synthesis.
Several recent works have extended diffusion-based grasp synthesis in various directions. Chen et al. [28] propose using stochastic interpolants to sample from an informed prior distribution instead of a Gaussian and bridge to the target distribution. Singh et al. [29] use a part-guided diffusion approach to generate grasps tailored to areas of interest instead of the whole object. Barad et al. [30] propose learning a latent space where diffusion on grasps is performed, and use language to condition on grasp regions. Carvalho et al. [31] propose a denoising diffusion probabilistic model for grasp synthesis, and optimize the grasps at test-time using collision spheres. Similarly, we propose a test-time grasp pose optimization technique that boosts performance without additional training on expert grasps, demonstrating controllable generation. Lim et al. [32] proposes a novel equivariant lifting layer for constructing -equivariant time-dependent velocity fields in flow matching.
@@ II-B Neural Shape Inference
Recent work has addressed the interdependence of grasp synthesis and geometry reconstruction. Breyer et al. [42] compute a TSDF volume of the scene, which is fed to a neural network to estimate grasp pose and quality at discrete locations on the 3D grid. Song et al. [43] propose to combine dense prediction with multimodal grasp generation, in an attempt to address the limitations of volumetric discretization. Jauhri et al. [44] pose grasping as a rendering problem by jointly learning to render a locality of the object along with the grasping functions in a shared feature space. Different from this direction of work, we consider shape learning as an implicit feature learning problem, to drive grasp synthesis. Here, the shape inference and grasp synthesis problems can be decoupled and learned in a more scalable manner than attempting to learn jointly using expert grasp data.
@@ III-B Signed Distance Fields
Signed Distance Fields (SDFs) implicitly represent geometry by providing, at every point in space, the distance to the nearest surface, with a sign denoting the interior (negative) or exterior (positive). Formally, we let denote the signed distance field. The surface is then represented as the zero level-set of its SDF: . Its gradient satisfies the Eikonal equation: . SDFs are differentiable almost everywhere, and where the gradient exists, its negative points toward the nearest surface. In recent years, much work has focused on implicitly representing 3D geometry via neural SDFs [26, 22], where the SDF is parameterized by a neural network. This allows querying at arbitrary resolutions with a small memory footprint and provides gradients via automatic differentiation. In this work, we use SDFs to learn shape features, which we transfer to stable grasp synthesis. We also use a learned SDF of the robot gripper for test-time pose optimization of the gripper given object point cloud.
@@ IV-A Shape Inference via Variational Autoencoding
High-quality grasp synthesis requires understanding geometric features that can transfer across object shapes and be robust to imperfect observations. While inferring shape features from point clouds is an extensively researched task [48, 38, 41], multimodal frameworks that learn shape priors in a latent space offer superior robustness to observation noise and capture a broader space of plausible shape hypotheses [49, 50]. Following Chou et al. [41], we learn SDF representations from object point clouds using a PointNet-based variational autoencoder (PointVAE) architecture that encodes useful geometric priors. Mapping (noisy) shape measurements to the shape latent space reduces the impact of imperfect shape measurements and enforces the effect of key shape features guided by geometric priors encoded in the shape latent space.
@@ IV-C Test-time Pose Optimization
where balance the influence of the objective functions with the score function. In our experiments, we use . We use the SIREN architecture [21] to model due to its infinite differentiability and proven effectiveness for smooth, differentiable objectives [27].
@@ V Experiments
We evaluate our approach against state-of-the-art multimodal grasp synthesis frameworks: cVAE [7], -Diff [8], Bridger [28], GLDM [30], and EGF [32]. All methods are trained on an NVIDIA A100 (40 GB) GPU for hours (72 hours for partial point clouds) using ten object categories from the ACRONYM dataset [58]: Book, Bottle, Bowl, Cap, CellPhone, Cup, Hammer, Mug, Scissors, and Shampoo. We implement our method in PyTorch [59] and evaluate grasp success and stability in IsaacGym [60]. Additionally, we demonstrate zero-shot sim-to-real transfer on a pick-and-place task without additional training.
@@ V-B Grasp Synthesis
[width=1.09]svg-inkscape/plot_sr_partial_box_svg-tex.pdf_tex
[width=1.09]svg-inkscape/plot_emd_partial_box_svg-tex.pdf_tex
[width=]svg-inkscape/plot_legend_partial_box_svg-tex.pdf_tex
@@
We evaluate on 52 unseen object instances from the ACRONYM dataset, measuring average success rate (SR) and an empirical Earth Mover’s Distance (EMD) with respect to the ground truth distribution as indicators of grasp quality and diversity, respectively. Fig.˜4 presents our main results.
@@
The cVAE baseline suffers due to training on a limited set of objects for grasp synthesis, while also demonstrating a lack of diversity due to mode collapse [8]. Furthermore, to boost its performance the cVAE depends on guidance from a grasp classifier [7] which is expensive to train. GLDM’s emphasis on learning a latent grasp space, which we hypothesize is not needed due to already working in a low dimension, proves less effective. Although the SMLD formulation of -Diff achieves reasonable performance, we believe it is limited by the VNN-based shape encoder. Bridger improves on -Diff by employing stochastic interpolants, achieving a higher SR with significantly lower inference time. EGF uses flow matching to achieve fast sampling, but generates less feasible grasps, which we attribute to requiring longer training times to handle the challenging set of objects. Our method achieves the highest success rate of , outperforming the previous state-of-the-art by , while maintaining competitive diversity (EMD = ) and reasonable inference time ( seconds).
@@
Key to our performance is an improved grasp synthesis architecture, a robust latent space for shape feature learning, and pre-training the shape backbone for object reconstruction. To see that the improved grasp synthesis architecture alone does not contribute to this performance improvement, we train "w/VNN" which yields lower performance. We also show the usefulness of pre-training for learning shape features by training "w/VNN+", where the VNN is pretrained via a surface reconstruction objective (described in Sec.˜V-A). Doing so boosts grasp synthesis performance compared to "w/VNN" but underperforms compared to Ours. This observation is supported by the shape inference results in Sec.˜V-A. Finally, our proposed test-time pose optimization method (Ours + GrOpt) boosts success rate to by encoding collision avoidance and pinch stability objectives. While this optimization slightly reduces diversity due to local refinement, it improves grasp quality and reduces variance. Qualitative results in Fig.˜6 illustrate the higher grasp quality achieved by our method across a diverse set of objects.
V-C Sparse/Partial Point Clouds
Real-world perception systems often produce sparse or noisy point clouds due to sensor limitations and environmental factors. We evaluate robustness by testing on point clouds of varying density, from 128 to 2048 points, where the training default is 1024. We exclude GLDM in this analysis as it does not ingest point clouds of varying sizes. Fig.˜7 shows that methods using standard point cloud encoders experience significant performance degradation as point cloud sparsity increases. In contrast, our approach has consistent performance across all sparsity levels, with success rates remaining above even with highly sparse inputs. This robustness stems from our variational shape inference approach, which learns a structured latent space that captures geometric priors and compensates for sparsity.
In many practical scenarios, using the full surface point cloud is not feasible. We evaluate our method’s efficacy in such scenarios by training on partial surface point clouds rendered from arbitrary camera view-points centered on the object. During training, we filter ground truth grasps for which the swept volume of the gripper does not contain any point from the rendered point cloud, i.e., only grasps that encapsulate the partial point cloud. Results are presented in Fig.˜5, where our method outperforms the baselines at success rate while abiding by the ground truth grasp distribution for the corresponding partial point cloud.
V-D Real-world Deployment
To validate the effectiveness of our variational shape inference approach with real sensor data, we deploy our method on a Franka Research 3 robot arm equipped with a Panda Hand gripper in a pick-and-place task. The experimental setup uses a Realsense D435i camera to capture the robot’s workspace, with segmented object point clouds extracted using SAM [63] as shown in Fig.˜8. Collision-free trajectories to and from the generated grasp poses are computed using the RRT-Connect implementation from VAMP [64]. Video demonstrations are provided in the supplementary material.
All methods are deployed without additional training or fine-tuning, evaluating zero-shot sim-to-real transfer capabilities. As presented in Tbl.˜II, our method achieves consistently high performance across all test objects, with particularly robust results for challenging geometries such as the Jug and Klein Bottle (K-Bottle). The Jug is significantly larger compared to other objects, and the Klein Bottle features smoothly varying curvature and complex geometry. Furthermore, real-world factors such as sensor noise and point cloud registration errors exacerbate the difficulty for methods that rely on standard point cloud encoders, while highlighting the advantage of our variational shape inference approach (trained only on noise-free point clouds).
VI Conclusion
In this work, we bring together approaches from generative modeling and implicit neural representations to enable robust and generalizable grasp synthesis. We employ a shape inference formulation parameterizing SDFs conditioned on object point clouds, and use the learned model to formulate grasp synthesis via a denoising diffusion model. Our proposal achieves superior performance compared to the state of the art in multimodal grasp synthesis, and demonstrates robustness in the presence of observation noise. Additionally, our test-time optimization technique further improves grasp quality without additional training on expert data.
This work furthers the case for the usefulness of transferring geometry understanding approaches to grasp synthesis [6]. The demonstrated robustness to point cloud sparsity and the successful real-world experiments underscore the practical value of variational shape inference for robotic manipulation. Therefore, in the future, large pretraining for shape inference may immediately be made useful for grasp synthesis in a few-shot manner [65].
A bottleneck in our presented approach is slower inference times. This is mainly due to the repeated automatic differentiation calls during Langevin Dynamics. However, we note that this does not pose a limitation to our shape inference formulation. As such our approach may be coupled with faster diffusion-based sampling techniques [28].